# V2X-R: Cooperative LiDAR-4D Radar Fusion with Denoising Diffusion for 3D Object Detection

### v3

@@ 2 Related Work
3D object detection has garnered significant attention, with mainstream approaches categorized as LiDAR-based [59, 18, 61, 40, 36, 37, 46, 3, 65, 47] or LiDAR-camera fusion-based [44, 43, 24, 58, 20]. The latter achieves superior performance by combining LiDAR’s depth and geometry information with the fine-grained details from images. However, LiDAR’s weather sensitivity leads to poor performance in adverse conditions, prompting research into weather-robust 3D object detection methods [15, 23, 8]. Some approaches utilize weather-robust 3D Radar [31, 38, 21, 22] or 4D radar [42, 41, 1, 16] fusion with LiDAR. 4D radar has received more attention due to higher resolution and elevation, but cooperative LiDAR-4D radar fusion remains unexplored.
@@ 3.2 Sensor configuration
As described in Table 1, each Connected Autonomous Vehicle (CAV) and Infrastructure is equipped with four cameras; a 64-channel LiDAR sensor featuring a detection range of 120 meters; a radar sensor with a vertical field of view of 30 degrees and a maximum detection range of 150 meters. The vehicle is equipped with a Global Navigation Satellite System (GNSS) with an altitude noise of 0.001 meters, while the roadside unit (RSU) has an altitude noise of 0.05 meters. Finally, based on the above configuration, our V2X-R contains a total of 12,079 scenarios with 37,727 frames of LiDAR and 4D radar point clouds, 150,908 images, and 170,859 annotated 3D vehicle bounding boxes.
@@
1x LiDAR 64 channels,120m range, -25°to 2° vertical FOV, 0.02 noise standard deviation, 20 Hz rotation frequency
1x 4D radar 150m range, 120° horizontal FOV, 30° vertical FOV
GPS & IMU Vehicle GNSS:altitude noise 0.001m; Vehicle IMU: heading noise 0.1°, speed noise 0.2m/s; RSU GNSS:altitude noise 0.05 m,
@@ 4.2 Fusion Pipeline
In the cooperative LiDAR-4D radar fusion pipeline, there are three types of agents: Ego Vehicle (E), Connected Automated Vehicle (C), and Infrastructure (I). Each agent collects LiDAR and 4D radar point cloud data, forming the multi-agent multi-modal input $\mathbf{X}=\{X^{L}_{C},X^{L}_{E},X^{L}_{I},X^{R}_{C},X^{R}_{E},X^{R}_{I}\}$. As illustrated in the Fig. 4(a), subsequent fusion consists of four stages:
@@ 4.3 Multi-modal Denoising Diffusion (MDD)
Therefore, addressing the dense noise challenge is crucial for weather-robust 3D object detection. Point-level semantic segmentation for denoising is time-consuming, struggles with intertwined noise and objects, and is incompatible with common multi-agent feature fusion strategies. On the other hand, simple feature-level denoising leads to poor fitting due to the complex distribution of weather noise in the feature. To overcome these limitations, we propose the Multi-modal Denoising Diffusion (MDD) for denoising noisy LiDAR features. As shown in Fig. 4(b), we resampled the complex weather noise distribution to a Gaussian distribution through the diffusion process, and then denoised it with the weather-robust 4D radar feature.
@@ 5.2 Benchmark Models
Cooperative LiDAR-based 3D object detectors. We pick existing method including AttFuse [54], V2XViT [53], CoBEVT [55], CoAlign [26], Where2comm [12], AdaFusion [17], SCOPE [60], MACP [27] and SICP [32] for cooperative LiDAR-based benchmarking analysis.
@@
V2XViT [53] ECCV2022 85.0/82.2/64.9 90.1/89.0/77.7
Attfuse [54] ICRA2022 86.0/82.2/66.9 91.4/89.6/80.2
Where2comm [12] NeurIPS2023 83.7/80.2/56.8 88.1/86.1/69.2
SCOPE [60] ICCV2023 76.0/74.7/60.9 81.4/72.9/67.0
CoBEVT [55] CoRL2023 87.6/84.8/71.0 92.3/89.4/82.5
CoAlign [26] ICRA2023 86.1/84.6/72.1 89.6/88.9/83.3
AdaFusion [17] WACV2023 88.1/86.9/75.6 92.7/90.6/84.8
SICP [32] IROS2024 81.1/77.6/58.1 84.6/82.2/66.7
MACP [27] WACV2024 72.8/70.9/60.0 83.7/83.1/75.5
@@
PFA-Net [50] ITSC2021 76.9/68.0/39.3 85.1/79.9/52.5
RTNH [29] NeurIPS2022 71.7/62.2/34.4 73.7/67.7/41.9
V2XViT [53] ECCV2022 71.1/64.3/31.1 80.9/73.8/42.7
AttFuse [54] ICRA2022 75.3/66.5/36.1 81.8/75.4/48.2
Where2comm [12] NeurIPS2023 69.4/62.1/26.6 77.8/72.9/41.5
SCOPE [60] ICCV2023 61.9/59.3/47.9 73.0/71.6/51.6
CoBEVT [55] CoRL2023 78.9/71.9/37.5 86.4/81.2/54.7
CoAlign [26] ICRA2023 65.8/59.2/34.7 76.9/70.2/46.2
AdaFusion [17] WACV2023 77.8/72.5/42.9 82.2/78.1/55.5
SICP [32] IROS2024 70.1/60.6/32.4 71.5/63.5/33.4
Table 3: Experimental 3D object detection results of various cooperative LiDAR-based methods on the validation and testing of our V2X-R dataset in different IoU (0.3,0.5,0.7). The representations of $\dagger$ and $\ddagger$ as described in 5.2.
@@
InterFusion [42] IROS2023 SA2MA 81.23 77.33 52.93 87.91 86.51 69.63
L4DR [16] Arxiv2024 SA2MA 84.58 82.75 70.29 90.78 89.62 82.91
V2XViT [53] ECCV2022 SM2MM 85.23 83.90 69.77 91.99 91.22 83.04
AttFuse [54] ICRA2022 SM2MM 86.14 84.30 70.72 92.20 90.70 84.60
SCOPE [60] ICCV2023 SM2MM 78.79 77.96 62.57 83.38 82.89 70.00
Where2comm [12] NeurIPS2023 SM2MM 87.62 85.58 69.61 92.20 91.00 82.04
CoBEVT [55] CoRL2023 SM2MM 86.69 85.72 75.93 94.00 93.23 86.68
CoAlign [26] ICRA2023 SM2MM 87.16 85.54 73.79 92.14 91.21 83.77
AdaFusion [17] WACV2023 SM2MM 89.24 87.31 74.86 93.10 92.15 85.48
SICP [32] IROS2024 SM2MM 82.46 79.44 61.51 86.19 84.20 68.15
@@
1.L4DR [16] L - - - - - - - - -
2.L4DR [16] L+4DR 78.88 75.87 59.10 85.00 80.09 61.48 90.66 89.29 80.42
3.L4DR w/MDD L+4DR 82.89 80.52 65.80 86.00 81.60 63.66 90.44 89.15 80.36
Improvement (3-2) +4.01 +4.65 +6.70 +1.00 +1.51 +2.18 -0.22 -0.14 -0.06
1.AttFuse [54] L 62.48 61.53 51.11 74.10 71.35 57.30 87.14 85.00 74.40
2.AttFuse [54] L+4DR 79.63 77.37 63.71 85.00 80.64 62.89 91.77 90.43 82.22
3.AttFuse w/MDD L+4DR 83.78 81.19 66.86 87.37 83.90 68.64 91.61 90.27 82.43
Improvement (2-1) +17.15 +15.84 +12.60 +10.90 +9.29 +5.59 +4.63 +5.43 +7.82
Improvement (3-2) +4.15 +3.82 +3.15 +2.37 +3.26 +5.73 -0.16 -0.10 +0.21
@@ 5.3 Benchmark Analysis
Cooperative LiDAR-based 3D object detectors. Table 2 depicts the performance of the selected cooperative LiDAR-based 3D object detectors. A large number of combined results from these models validate the reliability of our V2X-R dataset. It is worth noting that the performance of the newest methods (SICP [32], MACP [27]) is not as good as that of earlier methods (e.g. COBEVT [55], CoAlign [26], AdaFusion [17]). This is because these newest methods do not solely focus on mainstream performance: SICP emphasizes preserving the ability of individual perception, while MACP employs transfer fine-tuning instead of retraining to avoid the overhead.
Cooperative 4D radar-based 3D object detectors. We also pioneered the exploration of cooperative 4D radar-based methods on the V2X dataset. As shown in Table 3, the performance of 4D radar-based models is generally lower than that of LiDAR-based models in Table 2. This can be attributed to the significantly lower resolution of 4D radar than LiDAR, which is a limitation hindering the independent use of 4D radar in single-agent scenarios. However, we also observed that cooperative 4D radar-based methods can achieve the excellent performance of 86.4% 3D mAP@testing (IoU=0.3) (CoBEVT [55]). This shows that 4D radar can achieve independent perception through cooperative communication, proving the research significance of the cooperative 4D radar-based approach.
Cooperative LiDAR-4D radar fusion 3D object detectors. Finally, we implemented and evaluated a series of cooperative LiDAR-4D radar fusion methods following the methodology described in sec. 4.2. As shown in Table 4, by comparing with Tables 2 and 3, it can be observed that cooperative LiDAR-4D radar fusion achieves the best performance compared to other single-modality approaches. For example, V2XViT [53], AttFuse [54], and CoBEVT [55] exhibit 5.34%, 5.80%, and 4.18% 3D mAP@testing (IoU=0.7) improvements through LiDAR-4D radar fusion, respectively. These results convincingly validate that cooperative 4D radar can bring considerable performance enhancements to the LiDAR single-modality approach.
@@ 5.4 Multi-modal Diffusion Denoising Analysis
Performance comparison under different simulated weather on V2X-R dataset. We selected three models, AttFuse [54] and L4DR [16], which include SA2MA and SM2MM fusion strategies, to evaluate the performance under different weather conditions. As shown in Table 5, the LiDAR-4D radar fusion can achieve superior performance in various weather conditions. AttFuse achieved 17.17% 3D mAP@snow (IoU=0.3) improvement when compared to LiDAR-only (Improvement (2-1)), respectively. Moreover, our designed MDD module improved the basic LiDAR-4D radar fusion model performance by up to 5.73% (AttFuse, 3D mAP@fog) and 6.70% (L4DR, 3D mAP@snow) under adverse weather conditions. Meanwhile, our MDD module hardly affects the normal weather performance, which validates the effectiveness and reliability of our MDD module.
@@
Performance comparison under different real-world weather on K-Radar dataset. We further conducted experiments on the K-Radar single-agent real-world dataset. As shown in Table 6, our MDD module demonstrates significant performance improvements in real-world adverse weather, with 3D mAP gains of 5.20% and 5.97% for the Sedan and Bus classes, respectively. Consistent with the results on V2X-R dataset, MDD has rarely impacted the normal weather performance, resulting in overall improvements of 4.99% and 3.63%. These results validate the effectiveness of the MDD module in the presence of dense noise under adverse weather, whether single or multi-agent perception. Detailed evaluation metrics and performance of each weather can be found in the supplementary materials.
@@
Effect of each component. We systematically evaluated each component, with the results summarized in Table 7. We initially attempted to denoise LiDAR features using Vanilla-Unet (row 2st) but found the results unsatisfactory. The complex distribution of LiDAR noise in multi-agent settings is difficult to directly fit. We significantly improve performance by transforming the noise distribution into Gaussian distribution through a multi-step diffusion process (row 3rd). Finally, by incorporating weather-robust 4D Radar features as conditional information for multi-modal denoising (row 4th), the model achieves the best performance of 87.37/83.90/86.64 3DmAP@Fog.
Effect of diffusion layers and loss weight. Finally, we investigate the impact of diffusion layers $\mathcal{T}$ and the loss weight of MDD in our MDD module. Table 8 presents representative results. We first explore the MDD loss weight with $\mathcal{T}=2$, considering both constant weights and the epoch-varying weight $\gamma(e,\psi)$ in eqs. 9. The best performance is achieved with $\gamma(e,\psi=3)$ when $\mathcal{T}=2$. We further experiment with different values of $\mathcal{T}$. When $\mathcal{T}=2$ and the loss weight is $\gamma(e,\psi=3)$, our MDD module achieves the best performance of 87.37/83.90/86.64 3DmAP@Fog. Although MDD inevitably introduces an additional inference time of 32 ms, it significantly improves weather robustness and still maintains real-time (about 20 FPS). These results validate the effectiveness of our designed weight $\gamma(e,\psi)$ and real-time of our MDD module.
@@ 6 Conclusion and Discussion
Limitations and future work. While our work provides a strong foundation, deeper exploration of cooperative LiDAR and 4D radar fusion remains underdeveloped. A compelling research direction is the full utilization of multi-agent and multi-modal information for more robust 3D object detection.

### v4

@@ 2 Related Work
3D object detection has garnered significant attention, with mainstream approaches categorized as LiDAR-based [62, 19, 64, 43, 39, 40, 49, 3, 67, 50] or LiDAR-camera fusion-based [47, 46, 27, 61, 21]. The latter achieves superior performance by combining LiDAR’s depth and geometry information with the fine-grained details from images [26, 25, 17]. However, LiDAR’s weather sensitivity leads to poor performance in adverse conditions, prompting research into weather-robust 3D object detection methods [15, 24, 8]. Some approaches utilize weather-robust 3D Radar [34, 41, 22, 23] or 4D radar [45, 44, 1, 16] fusion with LiDAR. 4D radar has received more attention due to higher resolution and elevation, but cooperative LiDAR-4D radar fusion remains unexplored.
@@ 3.2 Sensor configuration
As described in Table 1, each Connected Autonomous Vehicle (CAV) and Infrastructure is equipped with four cameras; a 64-channel LiDAR sensor featuring a detection range of 120 meters; a radar sensor with a vertical field of view of 30 and a maximum detection range of 150 meters. The vehicle is equipped with a Global Navigation Satellite System (GNSS) with an altitude noise of 0.001 meters, while the roadside unit (RSU) has an altitude noise of 0.05 meters. Finally, based on the above configuration, our V2X-R contains a total of 12,079 scenarios with 37,727 frames of LiDAR and 4D radar point clouds, 150,908 images, and 170,859 annotated 3D vehicle bounding boxes.
@@
1x LiDAR 64 channels,120m range, -25to 2 vertical FOV, 0.02 noise standard deviation, 20 Hz rotation frequency
1x 4D radar 150m range, 120 horizontal FOV, 30 vertical FOV
GPS & IMU Vehicle GNSS:altitude noise 0.001m; Vehicle IMU: heading noise 0.1, speed noise 0.2m/s; RSU GNSS:altitude noise 0.05 m,
@@ 4.2 Fusion Pipeline
In the cooperative LiDAR-4D radar fusion pipeline, there are three types of agents: Ego Vehicle (E), Connected Automated Vehicle (C), and Infrastructure (I). Each agent collects LiDAR and 4D radar point cloud data, forming the multi-agent multi-modal input $\mathbf{X}=\{X^{L}_{C},X^{L}_{E},X^{L}_{I},X^{R}_{C},X^{R}_{E},X^{R}_{I}\}$. As illustrated in Fig. 4(a), subsequent fusion consists of four stages:
@@ 4.3 Multi-modal Denoising Diffusion (MDD)
Therefore, addressing the dense noise challenge is crucial for weather-robust 3D object detection. Traditional point-level segmentation for denoising is time-consuming and incompatible with common multi-agent feature fusion strategies. On the other hand, simple feature-level denoising leads to poor fitting due to the complex distribution of weather feature noise. To overcome these limitations, we propose the Multi-modal Denoising Diffusion (MDD) for denoising noisy LiDAR features. As shown in Fig. 4(b), we resampled the complex weather noise distribution to a Gaussian distribution through the diffusion process, and then denoised it with the weather-robust 4D radar feature.
@@ 5.2 Benchmark Models
Cooperative LiDAR-based 3D object detectors. We pick existing method including AttFuse [57], V2XViT [56], CoBEVT [58], CoAlign [29], Where2comm [12], AdaFusion [18], SCOPE [63], and SICP [35] for cooperative LiDAR-based benchmarking analysis.
@@
V2XViT [56] ECCV2022 83.47/80.65/63.48 89.44/88.40/77.13
Attfuse [57] ICRA2022 86.69/82.58/66.56 91.21/89.51/80.01
Where2comm [12] NeurIPS2023 85.31/82.65/64.35 85.59/84.27/73.13
SCOPE [63] ICCV2023 79.43/77.35/65.08 80.67/79.88/72.00
CoBEVT [58] CoRL2023 86.65/84.59/70.30 91.41/90.44/81.06
CoAlign [29] ICRA2023 84.43/82.29/70.68 88.12/86.99/80.05
AdaFusion[18] WACV2023 88.19/86.96/75.55 92.72/91.64/84.81
SICP [35] IROS2024 81.08/77.56/58.10 84.65/82.18/66.73
@@
PFA-Net [53] ITSC2021 75.45/66.66/38.30 84.93/79.71/52.46
RTNH [32] NeurIPS2022 72.00/62.54/34.65 73.61/67.63/41.86
V2XViT [56] ECCV2022 68.58/61.88/29.47 80.61/73.52/42.60
AttFuse [57] ICRA2022 71.83/63.33/34.15 81.34/74.98/47.96
Where2comm [12] NeurIPS2023 68.99/56.55/25.59 79.21/72.88/36.15
SCOPE[63] ICCV2023 71.97/67.03/53.23 72.20/69.12/53.41
CoBEVT [58] CoRL2023 73.48/66.82/34.48 85.74/80.64/54.34
CoAlign [29] ICRA2023 75.05/68.11/41.20 81.69/75.74/52.01
AdaFusion [18] WACV2023 75.60/70.33/41.11 81.95/77.84/55.32
SICP [35] IROS2024 70.83/62.79/34.82 71.94/65.17/36.44
Table 3: Experimental 3D object detection results of various cooperative 4D Radar-based methods on the validation and testing of our V2X-R dataset in different IoU (0.3,0.5,0.7). The representations of $\dagger$ and $\ddagger$ as described in 5.2.
@@
InterFusion [45] IROS2022 SA2MA 78.33 74.70 51.44 87.91 86.51 69.63
L4DR [16] AAAI2025 SA2MA 80.91 79.00 67.17 90.01 88.85 82.26
V2XViT [56] ECCV2022 SM2MM 85.43 83.32 66.23 91.21 90.07 79.87
AttFuse [57] ICRA2022 SM2MM 83.45 81.47 69.11 91.50 90.04 82.44
SCOPE [63] ICCV2023 SM2MM 79.84 79.10 65.31 84.14 83.61 73.15
Where2comm [12] NeurIPS2023 SM2MM 88.05 85.98 69.94 92.20 91.16 81.40
CoBEVT [58] CoRL2023 SM2MM 86.45 85.49 75.65 94.23 93.50 86.92
CoAlign [29] ICRA2023 SM2MM 87.08 85.44 73.66 91.13 90.19 83.73
AdaFusion [18] WACV2023 SM2MM 88.87 86.94 74.44 92.94 91.97 85.31
SICP [35] IROS2024 SM2MM 83.32 80.61 63.08 84.83 82.59 67.61
@@
L4DR [16] L+4DR 78.88 75.87 59.10 85.00 80.09 61.48 90.66 89.29 80.42
L4DR w/MDD L+4DR 82.89 80.52 65.80 86.00 81.60 63.66 90.44 89.15 80.36
Improvement +4.01 +4.65 +6.70 +1.00 +1.51 +2.18 -0.22 -0.14 -0.06
1.AttFuse [57] L 68.73 65.35 45.54 81.71 78.48 61.52 89.30 87.39 76.09
2.AttFuse [57] L+4DR 79.63 77.37 63.71 85.00 80.64 62.89 91.15 89.80 81.75
3.AttFuse w/MDD L+4DR 83.78 81.19 66.86 87.37 83.90 68.64 90.99 89.64 81.96
Improvement (2-1) +10.90 +12.02 +18.17 +3.29 +2.16 +1.37 +1.85 +2.41 +5.66
Improvement (3-2) +4.15 +3.82 +3.15 +2.37 +3.26 +5.73 -0.16 -0.16 +0.21
@@ 5.3 Benchmark Analysis
Cooperative LiDAR-based 3D object detectors. Table 2 depicts the performance of the selected cooperative LiDAR-based 3D object detectors. A large number of combined results from these models validate the reliability of our V2X-R dataset. It is worth noting that the performance of the newest methods (SICP [35]) is not as good as that of earlier methods (e.g. COBEVT [58], CoAlign [29], AdaFusion [18]). This is because SICP [35] not only focus on mainstream performance but also emphasize preserving the ability of individual perception.
Cooperative 4D radar-based 3D object detectors. We also pioneered the exploration of cooperative 4D radar-based methods on the V2X dataset. As shown in Table 3, the performance of 4D radar-based models is generally lower than that of LiDAR-based models in Table 2. This can be attributed to the significantly lower resolution of 4D radar than LiDAR, which is a limitation hindering the independent use of 4D radar in single-agent scenarios. However, we also observed that cooperative 4D radar-based methods can achieve the excellent performance of 85.7% 3D mAP@testing (IoU=0.3) (CoBEVT [58]). This shows that 4D radar can achieve independent perception through cooperative communication, proving the research significance of the cooperative 4D radar-based approach.
Cooperative LiDAR-4D radar fusion 3D object detectors. Finally, we implemented and evaluated a series of cooperative LiDAR-4D radar fusion methods following the methodology described in sec. 4.2. As shown in Table 4, by comparing with Tables 2 and 3, it can be observed that almost all cooperative LiDAR-4D radar fusion models achieve the best performance compared to other single-modality approaches. Especially, CoAlign [29], CoBEVT [58], and Where2comm [12] exhibit 3.68%, 5.86%, and 8.23% 3D mAP@testing (IoU=0.7) improvements through LiDAR-4D radar fusion, respectively. These results convincingly validate that cooperative 4D radar can bring considerable performance enhancements to the LiDAR single-modality approach. On the other hand, the table also shows that there is a significant difference in the benefits of adding 4DRadar across different methods, which indicates that exploring how to maximize the value of cooperative LiDAR-4D Radar fusion is still worth exploring.
@@ 5.4 Multi-modal Diffusion Denoising Analysis
Performance comparison under different simulated weather on V2X-R dataset. We selected three models, AttFuse [57] and L4DR [16], which include SA2MA and SM2MM fusion strategies, to evaluate the performance under different weather conditions. As shown in Table 5, the LiDAR-4D radar fusion can achieve superior performance in various weather conditions. AttFuse achieved 18.17% 3D mAP@snow (IoU=0.7) and 5.66% 3D mAP@normal (IoU=0.7) improvement when compared to LiDAR-only (Improvement (2-1)), respectively. Moreover, our designed MDD module improved the basic LiDAR-4D radar fusion model performance by up to 5.73% (AttFuse, 3D mAP@fog) and 6.70% (L4DR, 3D mAP@snow) under adverse weather conditions. Meanwhile, our MDD module hardly affects the normal weather performance, which validates the effectiveness and reliability of our MDD module.
@@
Performance comparison under different real-world weather on K-Radar dataset. We further conducted experiments on the K-Radar single-agent real-world dataset. As shown in Table 6, our MDD module demonstrates significant performance improvements in real-world adverse weather, with 3D mAP gains of 5.20% and 5.97% for the Sedan and Bus classes, respectively. MDD has rarely impacted the normal weather performance, resulting in overall improvements of 4.99% and 3.63%. These results validate the effectiveness of the MDD module in the presence of dense noise under real-world adverse weather.
@@
Effect of each component. We evaluated the effect of each component, as shown in Table 7. We initially attempted to denoise LiDAR features using Vanilla-Unet (row 2st) but found the results unsatisfactory. The complex distribution of LiDAR noise in multi-agent settings is difficult to directly fit. We significantly improve performance by transforming the noise distribution into Gaussian distribution through a multi-step diffusion process (row 3rd). Finally, by incorporating weather-robust 4D Radar features as conditional information for multi-modal denoising (row 4th), the model achieves the best performance of 87.37/83.90/86.64 3DmAP@Fog.
Effect of diffusion layers and loss weight. Finally, we investigate the impact of hyper-parameters in our MDD module, including number of diffusion layers and the MDD loss weight $\mathcal{T}$ (both constant weights and the epoch-varying weight $\gamma(e,\psi)$ in eqs. 9). As shown in Table 8, the best performance of 87.37/83.90/86.64 3DmAP@Fog is achieved with MDD loss weight $\gamma(e,\psi=3)$ and number of diffusion layers $\mathcal{T}=3$. Although MDD inevitably introduces an additional inference time of 32 ms, it significantly improves weather robustness and still maintains real-time (about 20 FPS). These results validate the effectiveness of our designed weight $\gamma(e,\psi)$ and real-time of our MDD module.
@@ 6 Conclusion and Discussion
Limitations and future work. While our work provides a foundation, deeper exploration of cooperative LiDAR and 4D radar fusion remains underdeveloped. A compelling research direction is the full utilization of multi-agent and multi-modal information for robust 3D object detection.
Acknowledgements. This work was supported in part by the National Natural Science Foundation of China (No.62171393), and the Fundamental Research Funds for the Central Universities (No.20720220064).